the difference between two positive integers is 50 and the ratio of these integers is 1 : 3 . find these integers

Answer by mananth(16946)

let the numbers be x & y (x>y)
x-y=50
x/y = 1/3
3x=y
3x-y=0
1 x + 1 y = 50.00 .............1
Family II
3 x -2 y = 0.00 .............2
Eliminate y
multiply (1)by 2
Multiply (2) by 1
2 x 2 y = 100.00
3 x -2 y = 0.00
Add the two equations
5 x = 100.00
/ 5
x = 20
plug value of x in (1)
1 x + 1 y = 50.00
20 + y = 50.00
y = 50.00 -20
y = 30.00
y = 30.00